OpenStudy (anonymous):

Please help!! If the cos of angle x is 8 over 17, and the triangle was dilated to be two times as big as the original, what would be the value of the cos of x for the dilated triangle? Hint: Use the slash symbol ( / ) to represent the fraction bar and enter the fraction with no spaces.

ganeshie8 (ganeshie8):

assume its a right triangle

ganeshie8 (ganeshie8):

after dilation, the sides will be 2 times longer : ` 2*8x`, `2*17x`

ganeshie8 (ganeshie8):

take the cos ratio again : cos = \(\dfrac{2*8x}{2*17x} = \dfrac{\not 2*8\not x}{\not2*17 \not x} = \dfrac{8}{17}\)

ganeshie8 (ganeshie8):

So, the ratio didnt change. looks the cos ratio is not affected by dilation ?

OpenStudy (anonymous):

it's supposed to be in fraction, so should I just put not effected?

ganeshie8 (ganeshie8):

Nope.

ganeshie8 (ganeshie8):

just put the old fraction

OpenStudy (anonymous):

ah! Ok! Thank you very much!

ganeshie8 (ganeshie8):

just put \(\dfrac{8}{17}\) okay ?

ganeshie8 (ganeshie8):

np.. u wlc :)